May 1 - Purchased office equipment for $30,000, paying $6,000 down and signing a 2-year, 12% (annual rate) note payable for the balance. The office equipment is expected to have a useful life of 10 years and a salvage value of $3,000. Straight-line depreciation is used.

Required: On the basis of the preceding information (both the transactions and additional information), prepare AJEs to adjust the company’s books as of December 31, 2020.
Answer #1
Date Account titles and explanation Debit Credit
2020
Dec 31. Depreciation expense (Note:1) 1800
Accumulated depreciation 1800
(Depreciation recorded)
Dec 31. Interest expense (Note:2) 1920
Interest payable 1920
(Interest accrued on note)
Note:1
Depreciation expense for the year=(Original cost-Salvage value)/Useful life=(30000-3000)/10=27000/10=$ 2700
But, Office equipment purchased on May 1
Hence,compute depreciation for 8 months(May to Dec)
Depreciation expense for 8 months=2700*(8/12)=$ 1800
Note:2
Note payable value=30000-6000=$ 24000
Interest expense for 8 months=24000*12%*(8/12)=$ 1920